OpenStudy (anonymous):

Pam bought several cartons of Giggle Fruit Drinks to sell during the track meet. She paid 50 cents per carton. She sold all but 15 of the cartons for $1 each. At the end of the day, her profit was $22.50. How many cartons of Giggle did she buy? @johnweldon1993

OpenStudy (johnweldon1993):

Since she bought them for 50 cents and sold them for a dollar, she made 50 cents profit on each one. If her total profit was $22.50, divide .50 into 22.50 to find that she sold 45 drinks. If she had 15 left over, she must have had 45 + 15 = 60 when she started. so she had bought 60 carton to begin with

OpenStudy (anonymous):

thts what i put and it didnt get it

OpenStudy (johnweldon1993):

hmm gimme a second then...

OpenStudy (johnweldon1993):

Oh duh...no its actually 75 cartons.... She paid 0.5x (.50 cents per carton) She sold all but 15 for $1 each, so she earned = 1(x -15) (x-15) - 0.5x = 22.50 x - 15 - 0.5x = 22.50 0.5x - 15 = 22.50 0.5x = 37.50 x = 75 So 75 cartons

OpenStudy (johnweldon1993):

And check it....75 cartons means she paid 37.50 for them all if she sold all but 15...then she sold 60 at 1 dollar each...so 60 if he profit was 22.50 then 60 - 37.50 = 22.50 Yes it does...so yeah 75 cartons